Find a sequence of three intergers with a common difference of three, such that three times the sum of the first two intergers minus five times the largest interger is 61.
Answer by edjones(8007)

Let the integers be x, x+3, x+6
.
3(x+x+3)-5(x+6)=61
6x+9-5x-30=61
x-21=61
x=82
.
Ed